OpenStudy (anonymous):

Find the slope and y-intercept of the graph of 5x-4y=20

OpenStudy (anonymous):

The equation for a line is y=mx+c So rearrange that formula to give that form

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

m is the gradient or the slope of the curve and to get the y intercept, you just make x=0

OpenStudy (anonymous):

how to find the value of m?

OpenStudy (anonymous):

So let's first rearrange the formula 5x-4y=20 4y=5x-20 y=5/4x-5

OpenStudy (anonymous):

So m=5/4

OpenStudy (anonymous):

is this okay? y=5/4x-5

OpenStudy (anonymous):

5x is supposed to be negative right?

OpenStudy (anonymous):

That's the equation for the line in the form y=mx+c So 5/4 is the slope

OpenStudy (anonymous):

how did the 20 become negative 5?

OpenStudy (anonymous):

You have to rearrange the formula into the form y=mx+c 5x-4y=20 -4y=-5x+20 4y=5x-20 Divide everything by 4 to make 1y instead of 4y y=5/4x-5

OpenStudy (anonymous):

\[y=\frac{ 5 }{ 4 }x-5?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks..

OpenStudy (anonymous):

Do you understand or do you need further explanation?

OpenStudy (anonymous):

you mean divide everything by -4?

OpenStudy (anonymous):

For -4y=-5x+20 Yes you do

OpenStudy (anonymous):

so the answer will be this way? \[y=\frac{ -5 }{ -4 } x+5\]

OpenStudy (anonymous):

No... When you divide a negative by a negative you will get a positive. So -5/-4 = 5/4 Also 20/-4 will give -5, not positive 5.

OpenStudy (anonymous):

You already solved the final answer which was \[y=\frac{ 5 }{ 4 }x-5\]

OpenStudy (anonymous):

oh.. yeah.. i see

OpenStudy (anonymous):

sorry..

OpenStudy (anonymous):

lol no problem you just need more practice so you can rearrange quicker

OpenStudy (anonymous):

can you help me with this too.? rite an equation of a line that has slope –5 and y-intercept 2.

OpenStudy (anonymous):

So the gradient of the line is -5, and when x=0 the y value = 2 y=mx+c (m is the gradient, c is the y intercept) y=-5x+2

OpenStudy (anonymous):

y=-5x+2 is the equation?

OpenStudy (anonymous):

thanks sir..

OpenStudy (anonymous):

Yes it is and no problem